A ball of mass 0.45 kg is fired with velocity 120 m/s into the barrel of a spring gun of mass 2.0 kg initially at rest on a frictionless surface. The ball sticks in the barrel at the point of maximum compression of the spring. No energy is lost to friction. What fraction of the ball's initial kinetic energy is stored in the spring?
Answer #1

All the KE of the ball is stored in the spring EXCEPT what goes into the KE of the ball+ gun.
Conservation of momentum: 0.45(120) = (2+0.45)Vfinal
So, Vfinal = 22.04 m/s and KEfinal = 0.5(2.0)22.04^2 = 485.797 J

Initial KE of the ball = 0.5(0.45)120^2 = 1296J
Energy stored in spring = 1296-485.797 = 810.20J

[Answer]